What is the A,G for the following reaction (in kJ mol-1) at 298 K? 2 Si (s) + 3 H2 (g) = SizH6 (g) The conditions for this reI keep getting 2.10 as the answer here but it is incorrect. Is there something I am doing wrong?
Answer #1

Answer: 250 g) + 34 (8) I sin 619) su = salsigh) - [axaa (si) + 3xA9°(H)] = 127.1936 – [o to] = 127.1936 kj/mol A9 = så + R7